how to find the equivalent wt of compounds in redox reactions sir?

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
equivalent weight of a compound in a redox reaction
=molecular wt/no of electron lost or gained during redox change by one molecule of that compound
for ex consider the reaction
Mn+4---------Mn+2
eq wt of Mn=at wt/change in oxdn number=55/2=27.5

FeC2O4-------Fe+3 +CO2
Fe+2 ------Fe+3 +e-
C2+3 --------2C+4 +2e-
FeC2O4----Fe+3 +2C+4 +3E-
eq wt of FeC2O4=M/3=144/3=48